### v1

# InfoSteer: Steering Information Utility in Language Model Post-Training

Abstract
Recent advancements in language models (LMs) gradually ushered in an era where post-training is crucial. Yet, post-training approaches such as supervised fine-tuning (SFT) do not guarantee effective use of knowledge acquired during pretraining. We therefore present InfoSteer, a lightweight method that encourages parametric information utilization in LMs during post-training. This is achieved via treating FFN layer as associate key-value memory, and promotes the use of stored memory vectors via forward-pass interventions or regularization during backpropagation. We find this simple guidance during post-training phase delivers consistent performance improvements across diverse model families–including Qwen, Gemma and Llama-spanning over 15 downstream tasks in both ID and OOD evaluations. Beyond performance gains, we also find that steered LMs can adaptively allocate information-placing more emphasis on generating semantically meaningful tokens, while using fewer resources on simple transition ones (e.g., ‘,’ or ‘and’). Our work underscores that vanilla post-training does not fully leverage pre-training potential, and steering LMs in latent representation space offers a promising approach that enhances both performance and interpretability.
1 Introduction
The contemporary training pipeline for LMs has standardized around a two-stage process: an initial pre-training phase on extensive, web-scale corpora, followed by a post-training phase utilizing smaller, more curated datasets (Ouyang et al., 2022; Touvron et al., 2023; Bai et al., 2023; Mesnard et al., 2024). A considerable body of literature suggests that the fundamental capabilities and knowledge of these models are predominantly instilled during the pre-training stage (Chung et al., 2022; Anil et al., 2023; Muennighoff et al., 2025). Subsequent post-training techniques are often understood as methods to better organize, elicit, or restructure these pre-existing abilities from the base model (Zhou et al., 2023; Rafailov et al., 2024; Guo et al., 2025; Swamy et al., 2025).
Despite this, an open question remains: do post-training methods actually encourage the model to fully utilize the information encoded during pre-training? In many cases, they may not-since models are not explicitly trained or incentivized to retrieve and apply such knowledge [^1] In this work, the terms ‘information’ and ‘knowledge’ are interchangeably used as descriptive language.optimally (Conmy et al., 2023; Chang et al., 2024; Du et al., 2024a). This insufficient use could result in suboptimal performance on downstream tasks, even when relevant knowledge is already stored in the model’s parameters (Bietti et al., 2023; Kim et al., 2025).
We therefore introduce InfoSteer, a prototype designed to encourage more efficient use of a model’s pre-trained knowledge during the post-training phase (Figure 1). Our approach draws on an associative memory view of the Transformer’s feed-forward network (FFN) layers (Geva et al., 2021; Meng et al., 2022a, b; Geva et al., 2023), wherein the first FFN layer acts as a content-dependent key and the second FFN layer as a learned value memory (key-value memory). Consequently, each layer’s output can be viewed as a weighted sum over key-value memory pairs, where each key acts as a coefficient controlling its corresponding memory.
InfoSteer controls the distribution of keys to manage the information-dense memory vectors learned from pretraining. The goal is to promote a high-entropy key distribution, enabling their corresponding memory vectors to be actively engaged during post-training. Specifically, we use two-sided strategies: (i) intervening on memory vectors with low corresponding key values during the forward pass, and (ii) regularizing the entropy of keys in the gradient flow during backpropagation. These methods can be seamlessly incorporated into a vanilla SFT pipeline for information-steered SFT.
Figure 1: Overview of our proposed InfoSteer framework. The interpretation of Transformer FFNs as key-value memory was introduced by Geva et al. (2021), and we provide further details in §3.
Our empirical evaluation demonstrates that InfoSteer yields consistent and notable performance enhancements across a variety of models—including Llama (Touvron et al., 2023), Qwen (Bai et al., 2023), and Gemma (Mesnard et al., 2024) at different scales. These improvements are demonstrated across a comprehensive suite of over 15 downstream tasks, encompassing both in-distribution (ID) and out-of-distribution (OOD) scenarios. Beyond these performance gains, information-steered LMs exhibit an adaptive information allocation strategy during token generation. Specifically, they dedicate more representational capacity to semantically rich and challenging tokens, while expending less on simpler, transitional tokens, indicating a more nuanced use of parametric information.
2 Related Work
Model Steering.
Model steering represents an emerging paradigm that aims to guide model behavior, focusing on controlled interventions in the model’s internal representation space (Subramani et al., 2022; Zou et al., 2023; Turner et al., 2024; Li et al., 2024; Cyberey and Evans, 2025). Many work has demonstrated the effectiveness of steering methods across various dimensions, including disentangling human-interpretable concepts (Rumelhart et al., 1986), such as linguistic features (e.g., gender, number) (Hewitt and Manning, 2019; Lasri et al., 2022; Wang et al., 2022; Hanna et al., 2023; Huang et al., 2024; Chang et al., 2025) and logical reasoning (Wu et al., 2024b; Xie et al., 2025). These approaches typically operate by introducing auxiliary objectives, modifying activation patterns, or implementing controlled perturbations during forward or backward passes (Rimsky et al., 2023; Scalena et al., 2024; Stolfo et al., 2024; Luo et al., 2025; Gur-Arieh et al., 2025; Bartoszcze et al., 2025). The key advantage of model steering lies in its ability to leverage the rich knowledge already encoded in pretrained weights while directing how this information is accessed and applied to downstream tasks (Geiger et al., 2023; Lee et al., 2024; Siddique et al., 2025; Soo et al., 2025; Deng et al., 2025). Our information-steered approach extends this paradigm by targeting the key distribution in transformer FFN layers, which influences which memory vectors are activated during computation.
Parametric Information.
From an information-theoretic perspective, quantifying the information encoded in a LMs’ high-dimensional parameters space remains a challenging problem (Achille et al., 2020; Bernstein and Yue, 2021). This challenge is closely tied to the ongoing discourse on parametric versus non-parametric knowledge storage in LMs (Tay et al., 2022; Ferrando et al., 2022, 2023; Xie et al., 2023; Deng et al., 2024; Ferrando and Voita, 2024; Du et al., 2024b) (or token-mixing in attn layer vs. channel mixing in FFN layer). The concept of channel mixing stems from the perspective of viewing the FFN layer as a key-value memory structure (Weston et al., 2014; Geva et al., 2021; Qiu et al., 2024; Kim et al., 2025) . Empirical evidence supports this interpretation, showing that individual neurons in FFN layers activate in response to specific semantic concepts and linguistic patterns (Dai et al., 2021; Liu et al., 2023; Niu et al., 2024). Furthermore, intervention studies have demonstrated that targeted modifications to these memory vectors can directly influence model outputs on knowledge-intensive tasks (Meng et al., 2022a, b; Geva et al., 2023; Hase et al., 2023; Yao et al., 2024; Wang et al., 2024), suggesting that factual information is strongly correlated with in this component.
3 Transformer Feed-Forward Layers as Key-Value Memories
We consider a standard autoregressive Transformer (Vaswani et al., 2017), which models the conditional probability of a sequence $x=(x_{1},\ldots,x_{T})$ as
$p(x)=\prod_{t=1}^{T}p(x_{t}\mid x_{<t}),$ (1)
using a stack of $L$ Transformer decoder blocks. Each block consists of two primary components: a masked self-attention layer and a position-wise feed-forward network (FFN).
In standard Transformer architectures, the position-wise FFN in each decoder block is given by:
$\text{FFN}(X)=\sigma(XW_{up}+b_{1})W_{down}+b_{2},$ (2)
where $X\in\mathbb{R}^{T\times d}$ is the input representation matrix for FFN, $W_{up}\in\mathbb{R}^{d\times d_{m}}$ and $W_{down}\in\mathbb{R}^{d_{m}\times d}$ are the up-projection and down-projection matrix respectively, and $\sigma$ is a nonlinearity activation function. Bias term $b_{1}$ and $b_{2}$ are included in the standard FFN formulation but omitted in the following equations for simplicity.
Let $h\in\mathbb{R}^{d}$ represent the embedding as a single token representation from the input matrix $X$. As illustrated in Figure 2, this structure can be interpreted as a soft key-value memory (Weston et al., 2014; Geva et al., 2021), where the intermediate activation $\sigma(hW_{\text{up}})\in\mathbb{R}^{d_{m}}$ defines the soft addressing weights, and each row of $W_{\text{down}}$ serves as a value vector. The FFN output of the input $h$ can be written as a weighted sum over value rows:
$\text{FFN}(h)=\sum_{i=1}^{d_{m}}\underbrace{\sigma(hW_{\text{up}})_{i}}_{\text{key coefficient}}\cdot\underbrace{(W_{\text{down}})_{i,:}}_{\text{value vector}}=k_{1}v_{1}+k_{2}v_{2}+\cdots+k_{d_{m}}v_{d_{m}}$ (3)
where each $k_{i}\in\mathbb{R}$ is a scalar key coefficient, and each $v_{i}\in\mathbb{R}^{d}$ is a value vector (memory vector). The key coefficients control the magnitude of the memory vectors’ contribution to the final prediction.
4 InfoSteer
LMs are commonly interpreted to store parametric knowledge within their dense layers, particularly in the FFN blocks. (Meng et al., 2022a; Geva et al., 2023; Kim et al., 2025). However, during post-training, there is no explicit guidance to encourage the model to utilize this knowledge for new alignment or retrieval tasks. InfoSteer is proposed to bridge this gap.
4.1 Key Motivation & Desierata
To steer the information utility of pretrained LMs, we aim to enhance the engagement of memory vectors during post-training. The core idea is to control the distribution of key coefficients in the position-wise FFN (see Figure 2). Specifically, we observe a dominant property of associative memory: if a key coefficient $k_{i}$ is significantly larger than another key coefficient $k_{j}$, the final prediction predominantly relies on the corresponding memory vector $v_{i}$, while $v_{j}$ is underutilized or overlooked. Our goal is to achieve two objectives:
  1. 1.
    Minimize the language modeling loss: Ensure the FFN output $\text{FFN}(h)=\sum_{i=1}^{d_{m}}k_{i}v_{i}$ contributes to accurate predictions.
  2. 2.
    Maximize the entropy of the key distribution: Encourage diverse engagement of memory vectors by promoting a higher-entropy key coefficient distribution.
Promoting entropy weaken the dominance of individual coefficients, encouraging more balanced activation of memory vectors $v_{i}$ and thus enabling richer parametric knowledge retrieval.
4.2 Methods: Intervention and Regularization
Figure 2: Key Design of InfoSteer. (a) An illustration of viewing the Transformer FFN as a key-value memory. The key acts as a control mechanism that determines the extent to which each memory vector is engaged. (b) and (c) Two general methods used to modulate the distribution of key coefficients, thereby encouraging the engagement of memory vectors during post-training.
Formally, let $\mathbf{k}^{(l)}=[k_{1}^{(l)},k_{2}^{(l)},\dots,k_{d_{m}}^{(l)}]\in\mathbb{R}^{d_{m}}$ denote the key coefficients at layer $l$, where $k_{i}=\sigma(hW_{\text{up}})_{i}$ and $W_{\text{up}}\in\mathbb{R}^{d\times d_{m}}$ is the up-projection matrix. We propose two complementary methods to increase memory vector engagement: intervention and regularization.
Intervention Method.
The intervention method directly modifies key coefficients to promote broader memory vector utilization. For each layer $l$, we identify the $p\%$ of key coefficients with the smallest value and adjust them to a value proportional to the layer’s average key coefficient.
Formally, let $\mathcal{I}^{(l)}$ denote the set of indices corresponding to the $p\%$ of elements in $\mathbf{k}^{(l)}$ with smallest value. For a hyperparameter $\alpha>0$, we update:
$k_{i}^{(l)}\leftarrow\alpha\cdot\frac{1}{d_{m}}\sum_{i=1}^{d_{m}}k_{i}^{(l)},\text{for}\enspace i\in\mathcal{I}^{(l)}.$ (4)
This adjustment ensures that previously minor coefficients contribute more significantly to the FFN output $\text{FFN}(h)=\sum_{i=1}^{d_{m}}k_{i}v_{i}$. $\alpha$ is controlling scalar to set the magnitude of steering.
Regularization Method.
The regularization method encourages a uniform key coefficient distribution by adding an entropy term to the loss function. Higher entropy reduces the dominance of individual coefficients, engaging more memory vectors. The modified loss function, across $L$ layers, is:
$\mathcal{L}=\mathcal{L}_{\text{LM}}-\lambda\sum_{l=1}^{L}H(\hat{\mathbf{k}}^{(l)}),$ (5)
where $\hat{\mathbf{k}}^{(l)}=\frac{\mathbf{k}^{(l)}}{\sum_{i=1}^{d_{m}}k_{i}^{(l)}}$ is the normalized version of original key distribution $\mathbf{k}^{(l)}$, $\mathcal{L}_{\text{LM}}$ is the language modeling loss, and $\lambda>0$ controls the regularization strength.
Both methods modulate the key coefficient distribution to improve the model’s ability to leverage parametric knowledge during post-training tasks.
4.3 MLP Variants
When adapting methods for MLP variants like SwiGLU, used in models such as Qwen (Bai et al., 2023), LLaMA (Touvron et al., 2023), and Gemma (Mesnard et al., 2024), we focus on the input to the down projection. SwiGLU modifies the standard FFN as:
$\text{SwiGLU}(h)=(\sigma(hW_{\text{gate}})\odot(hW_{\text{up}}))W_{\text{down}},$ (6)
where $\sigma$ is the silu activation, and $\odot$ denotes the elment-wise multiplication. Similar to standard two-layer FFN, we modulate the key coefficient distribution $\mathbf{k}=\sigma(hW_{\text{gate}})\odot(hW_{\text{up}})$.
5 Experiment
To evaluate the overall effectiveness of steered SFT against vanilla SFT, we conduct a comprehensive study across various downstream tasks and model sizes. Our primary focus is on achieving a holistic understanding of the overall model behavior after applying our methods, with particular attention paid to generalization performance. The details are provided below [^1] In this paper, we primarily focus on a generic method. But we also provide fine-grained steering methods in Appendix B and the ablation study of different layers in Appendix C..
5.1 Experiment Setup
Base Models.
We evaluated our methods on language models of varying sizes across three series: Qwen-2.5-1.5B and Qwen-2.5-7B for the Qwen series (Bai et al., 2023), LLaMA-3.2-1B and LLaMA-3-8B for the LLaMA (Touvron et al., 2023) series, and Gemma-2-2B and Gemma-2-9B for the Gemma (Mesnard et al., 2024) series.
Datasets.
We evaluate performance under two settings: in-distribution (ID) and out-of-distribution (OOD) evaluations.
In the ID setting, we consider a diverse set of downstream tasks ranging from knowledge-intensive to reasoning-intensive. These include BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2019), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2019), GSM8K (Cobbe et al., 2021), ARC-e, ARC-c (Clark et al., 2018), and OBQA (Mihaylov et al., 2018). No chain-of-thought (CoT) (Wei et al., 2023) rationales are provided.
For OOD setting, we primarly trained GSM8K and eval on other arithmetic datasets, including AddSub (Hosseini et al., 2014), SingleEQ (Koncel-Kedziorski et al., 2015), MultiArith (Roy and Roth, 2016), AQuA (Ling et al., 2017), MAWPS (Koncel-Kedziorski et al., 2016), and SVAMP (Patel et al., 2021). In these benchmarks, chain-of-thought (CoT) rationales are typically included before the final answer. For all benchmarks, we use the same prompt templates as in Hu et al. (2023); Wu et al. (2024a). We also remove any leading and trailing whitespace from the dataset.
Baseline.
We use standard SFT as our baseline. For information-steered SFT, we evaluate intervention and regularization methods, which share the goal of enhancing memory engagement during training. The default hyperparameters for the intervention are set to $p$= 0.01 and $\alpha=1$, with $\lambda=0.01$ for entropy regularization. Other training details are provided in Appendix A.
5.2 General Performance Comparison
We use the default hyperparam setting as reported in § 5.1 to eval general performance difference. Table 1 presents the accuracy across nine commonsense reasoning datasets for Qwen, LLaMA, and Gemma models under various different steering strategies. We categorize the models into small-scale (1-2B parameters) and large-scale (7-9B parameters) groups.
Benefit of Model Steering.
Our experimental results, as presented in Table 1, provide compelling evidence for the efficacy of model steering techniques across different model architectures and parameter scales. Both proposed steering methods-steered SFT with intervention and steered SFT with regularization-consistently outperform their respective base models and vanilla SFT counterparts across all benchmarks. This pattern holds across all model families, demonstrating the robust, architecture-agnostic nature of our steering approaches.
Insufficient Utilization of Pretrained Knowledge in Modern LLMs.
The benefits of steering reveal a critical limitation in contemporary large language models: they substantially underutilize the knowledge acquired during pretraining when applied to downstream tasks. This inefficiency is strikingly evident across all three widely-used model families examined-Gemma, LLaMA, and Qwen. Specifically, the Gemma-2-9B base model achieves $90.1$% on HellaSwag, yet reaches $95.7$% with intervention steering-a $5.6$% improvement without additional pretraining or parameter increase. Similarly, LLaMA-3-8B shows a remarkable $5.5$% gain on HellaSwag with intervention steering over the base model. This pattern repeats across all architectures and benchmarks, suggesting that modern LLMs possess substantially more capabilities than they successfully deploy in standard fine-tuning regimes.
Table 1: Performance comparison of Qwen, LLaMA, and Gemma models with different training methods on eight datasets. To highlight improvements, we use blue for significant gains and green for moderate ones. All results are reported as the average scores over three independent runs.
Model Training Accuracy ($\uparrow$) BoolQ PIQA SIQA HellaS. WinoG. ARC-e ARC-c OBQA Small-Scale Models (1-2B parameters) Qwen-2.5-1.5B base model 64.2 78.5 74.3 80.1 76.4 76.9 61.2 75.8      + vanilla SFT 68.5 82.9 79.6 84.8 80.8 81.4 65.8 81.0      + steered SFT w. intervention 69.3 84.4 80.3 93.1 84.2 83.2 68.2 78.9      + steered SFT w. regularization 68.7 83.9 79.8 92.4 83.7 82.5 67.5 78.1 LLaMA-3.2-1B base model 65.6 75.3 74.2 78.9 77.8 74.5 60.1 76.3      + vanilla SFT 69.8 79.9 79.5 83.6 82.6 79.8 64.7 81.0      + steered SFT w. intervention 71.8 83.7 76.0 89.1 82.6 83.7 68.2 82.4      + steered SFT w. regularization 71.0 82.9 75.2 88.3 81.9 82.6 67.4 81.7 Gemma-2-2B base model 66.5 79.1 73.8 82.7 78.9 77.4 63.8 74.9      + vanilla SFT 70.2 83.4 78.1 87.5 83.3 82.7 68.4 80.1      + steered SFT w. intervention 72.5 85.6 79.3 90.2 85.8 85.3 71.9 83.7      + steered SFT w. regularization 71.8 84.9 78.5 89.3 85.0 84.7 71.0 82.9 Large-Scale Models (7-9B parameters) Qwen-2.5-7B base model 68.9 81.2 77.6 87.5 80.3 79.8 65.1 77.6      + vanilla SFT 72.4 84.9 81.5 92.4 84.2 84.2 69.6 82.8      + steered SFT w. intervention 74.1 86.3 81.8 95.1 87.2 86.2 73.7 84.2      + steered SFT w. regularization 76.4 85.7 81.0 94.3 86.5 85.4 72.9 83.4 LLaMA-3-8B base model 70.3 85.6 75.7 90.8 81.9 86.2 75.3 80.5      + vanilla SFT 74.6 89.3 79.9 93.5 85.6 90.5 80.4 85.8      + steered SFT w. intervention 77.1 90.2 82.0 96.3 87.4 92.4 81.6 87.5      + steered SFT w. regularization 76.5 89.5 81.2 95.6 86.8 91.7 80.9 86.8 Gemma-2-9B base model 71.6 86.3 77.2 90.1 82.5 87.5 77.8 81.7      + vanilla SFT 74.3 90.1 81.7 94.8 86.9 91.7 82.0 86.4      + steered SFT w. intervention 77.2 91.8 83.1 95.7 88.5 93.5 83.4 88.2      + steered SFT w. regularization 76.5 90.9 82.4 94.9 87.8 92.8 82.7 87.5
5.3 Ablations of Steering Magnitude
Table 2: Results under varying steering magnitudes. $p\%$ determines the proportion of keys being intervened, $\alpha$ controls the intervention strength, and $\lambda$ sets the regularization strength for the entropy.
Model Average Acc Base Model 71.4      + Vanilla SFT 72.6      + steered SFT w/ interv. ($p=1$, $\alpha=1$) 73.8      + steered SFT w/ interv ($p=1$, $\alpha=2$) 75.5      + steered SFT w/ interv ($p=2$, $\alpha=5$) 72.8      + steered SFT w/ reg. ($\lambda=-0.01$) 72.3      + steered SFT w/ reg. ($\lambda=0.01$) 73.4      + steered SFT w/ reg. ($\lambda=0.05$) 74.7
Table 2 summarizes the performance under various steering magnitudes and strategies. Incorporating our proposed steered SFT approach with intervention consistently enhances performance, with the best result of 75.5% achieved at intervention parameters $p=1$, $\alpha=2$. And excessive intervention magnitude may lead to suboptimal performance (72.8%), suggesting a balance is necessary for optimal results. Additionally, entropy-based regularization ($\lambda$) also demonstrates effectiveness in steering model performance. Positive entropy regularization ($\lambda=0.05$) significantly boosts accuracy to 74.7%, whereas negative regularization ($\lambda=-0.01$), which actively discourages information utility, leads to degraded performance, resulting in an accuracy of 72.3%.
5.4 Effectiveness of Model Steering Across Task Types
Table 3 presents performance improvements across five task types when applying different model steering strategies.
Table 3: Gain of Model Steering w/ Different Task Type.
Model Reading Comp. Knowledge Commonsense Reasoning Math Linguistic Base Model 72.3 70.1 65.4 63.7 78.2      + Vanilla SFT 73.8 70.6 66.1 65.7 77.9      + Steered SFT w/ interv. 78.1 74.9 68.5 66.8 78.0      + steered SFT w/ reg. 77.4 73.4 67.9 66.1 77.3 Average $\Delta$ +3.9 (1) +3.3 (2) +2.3 (3) +1.1 (4) -0.3 (5)
Both steered SFT variants outperform the base model and vanilla SFT. On average, these methods yield a $+3.9$ improvement in reading comprehension, $+2.3$ in commonsense reasoning, $+1.1$ in math, and $+3.3$ in knowledge tasks. However, linguistic tasks show a slight performance drop ($-0.3$). These results highlight the effectiveness of targeted model steering, particularly for knowledge-intensive and reasoning-heavy tasks, while linguistic tasks appear to benefit less from encouraging increased parametric knowledge utilization.
5.5 Out-of-Distribution Evaluation
Table 4: Results for ID and OOD evals. The OOD performance is reported as average score across five benchmarks.
Model ID Eval OOD Eval Base Model 63.7 85.3      + Vanilla SFT 65.7 (+2.0) 83.7 (-1.6)      + Steered SFT w/ interv. 66.8 (+3.1) 86.6 (+1.3)      + steered SFT w/ reg. 66.1 (+2.4) 86.0 (+0.7)
We trained on the GSM8K dataset to assess in-distribution (ID) performance and are evaluated on five arithmetic out-of-distribution (OOD) datasets-AddSub, MAWPS, MultiArith, SingleEq, and SVAMP-to evaluate generalization. This setup enables a comprehensive study of how different fine-tuning strategies transfer to novel problem structures.
Vanilla SFT Improves ID but Harms OOD.
As shown in Table 4, Vanilla SFT improves ID accuracy from 63.7 to 65.7, demonstrating its ability to fit the training distribution. However, this comes at the cost of generalization: its OOD performance drops from 85.3 to 83.7. This suggests that naive SFT may lead to overfitting and reduce robustness on unseen arithmetic tasks.
Steered SFT Enhances Both ID and OOD Performance.
Figure 3: OOD Model Performance Comparison across different mathematical reasoning datasets. Results are reported w/ average score of three separate runs.
In contrast, our proposed Information-steered SFT improve both ID and OOD scores. Specifically, the intervention-based steering reaches the highest ID score of 66.8 and the best OOD score of 86.6 . The regularization-based steering also yields consistent improvements (66.1 ID, +2.4; 86.0 OOD, +0.7). These results highlight that controlled interventions and regularizations during post-training can enhance generalization for LMs.
Fine-Grained Analysis Across OOD Datasets.
Figure 3 further illustrates performance across individual OOD datasets. Notably, Steered SFT w/ intervention achieves the best or comparable performance in nearly all settings. On the most challenging dataset, SVAMP, Vanilla SFT performs the worst, while both steered methods significantly improve accuracy. This demonstrates the robustness of InfoSteer on structurally diverse arithmetic problems.
6 Analysis
In this section, we address two key questions. First, compared to vanilla SFT, what kind of distribution shift does our method introduce in the key-coefficient distribution? Second, beyond improvements in performance, what other changes can be observed in the model’s behavior? These analyses provide deeper insights into both the effectiveness and interpretability of our steering strategy.
6.1 Distribution Shift after Steering
As shown in Figure 4, fine-tuning via standard SFT increases the number of activations in the low-key region, diverging from the distribution observed in the base model. This shift indicates that vanilla SFT encourages the model to rely on fewer memory vectors compared to its base model, potentially increasing the risk of overfitting to downstream tasks (as overfitting is observed in our previous experiment in § 5.5).
Figure 4: Key Coefficient Distribution Shift from Base Model. The gray line serve as baseline for the number of key in corresponding region.
Compared to vanilla SFT, which predominantly boosts activations in the low-key region, both steering methods adjust the distribution in more structured and nuanced ways. Regularization-based steering softens key allocation by reducing activations at the extremes—both low and high—while moderately increasing the usage of medium-range keys. In contrast, intervention-based steering results in a sharper redistribution, actively suppressing low-region keys and promoting greater use of medium and high-region keys. This targeted reshaping suggests that intervention-based steering, in particular, fosters more efficient and selective memory utilization within the model.
6.2 Self-Steering Leads to Better Interpretablity
In addition to performance improvements, we’re intrigued by how InfoSteer can enhance model interpretability. Recently, there has been growing interest in understanding the mechanisms to dissect the inner workings of LMs. For example, approaches like sparse autoencoders (SAEs) (Cunningham et al., 2023) involve training external NN to identify co-activation patterns in the model’s latent space. While this line of research is promising, we aim to explore a different perspective: that LMs might be more self-interpretable via proper guidance without extra training.
To investigate this, we first performed instruction tuning on Qwen-2.5-7B using UltraFeedback (Cui et al., 2023), a dataset comprising 6.4K high-quality instruction-following examples specifically designed to align model outputs with human-preferred instruction styles. Our goal was to analyze the model’s behavior after alignment, particularly under steering interventions. To this end, we computed token-level entropy over the key coefficient distributions as a proxy for how many memory vectors are activated during token generation. We refer to this metric as the Information Flux (IF) score, which quantifies the amount of information the model needs to use to generate each token.
Table 5 presents qualitative examples of token-level IF score across four prompt types. We observe meaningful patterns in both the highlighted (high-IF) and underutilized (low-IF) tokens.
Highlight Tokens Reflect Instruction Semantics.
The highlighted tokens in each response align closely with the semantic core of the instruction:
  • For Explain Concepts, the model concentrates high IF on the definitional span—“quantum computing is a type of computing that uses quantum mechanical phenomena...”—which directly addresses the instruction.
  • In Creative Writing, poetic and original terms such as “circuits,” “code,” and “digital symphony” are highlighted. These tokens carry the stylistic and imaginative weight of the response, indicating the model recognizes them as content-bearing.
  • For Arithmetic Calculation, the most salient tokens are in the initial planning phrase—“To calculate (31 $\times$ 31), you can use the formula...”—which frames the reasoning path. The mathematical expansion that follows shows diminished IF, likely due to repetition and procedural predictability.
  • In Game Strategy Reasoning, tokens like “winning strategy,” “losing position,” and “concept of winning and losing” are emphasized. These convey the strategic logic essential for solving the problem, demonstrating that the model assigns higher utility to abstract reasoning components.
Underutilized Tokens Encode Structure or Copy.
Across settings, transition tokens (e.g., “,”, “to”) consistently exhibit low IF, indicating minimal semantic contribution. Interestingly, math expressions directly reused from the prompt (e.g., “31 $\times$ 31”) are also under-highlighted, suggesting the model de-emphasizes copied content in favor of novel reasoning or planning segments. This pattern extends to common function words and syntactic markers that provide structure rather than content. For example, determiners (“the”, “a”), conjunctions (“and”, “or”), and pronouns (“it”, “they”) show consistently lower IF scores across all prompt types.
Table 5: Token-wise Information Utility (IF) across prompts. Highlight colors represent the IF intensity - Low,  medium,  High correspondingly.
Instruction Type Instruction Response (Token-level IF) Explain Concepts Explain the concept of quantum computing in simple terms. Quantum computing is a type of computing that uses quantum mechanical phenomena , such as superposition and entanglement , to perform operations on data . Creative Writing Write a short poem about artificial intelligence. In circuits and code we rise, A digital symphony of eyes. Our logic and data entwined, Arithmetic Calculation What is 31 $\times$ 31? To calculate ( 31 $\times$ 31 ), you can use the formula for squaring a binomial , which is (a+b) 2 = a 2 + 2ab + b 2 . Here, a = 30 and b = 1. So, we have: (30+1) 2 = ${30^{2}}$ + 2 $\times$ 30 $\times$ 1 + ${1^{2}}$ Game Strategy Reasoning (AIME2024) Alice and Bob play the following game. A stack of $n$ tokens lies before them. $\cdots$ Whoever removes the last token wins. Find the number of positive integers $n\leq 2024$ such that Bob always wins. To determine the number of positive integers ( n \ le q 2024 ) for which Bob has a winning strategy , we need to analyze the game using the concept of winning and losing positions . A position is called a losing position if the player whose turn
7 Conclusion
We introduce a lightweight and effective method for post-training that enhances parametric knowledge utilization in language models by steering the key-value dynamics in FFN layers. Our findings reveal a critical insight: modern LLMs substantially underutilize the knowledge acquired during pretraining when applied to downstream tasks, leaving significant performance potential unrealized. Through simple forward interventions and entropy-based regularization, steered SFT consistently improves both ID and OOD performance across diverse models and tasks. Beyond accuracy gains, our method encourages adaptive memory allocation and reveals interpretable information usage patterns, offering new insights into the internal behavior of post-trained LMs. These findings suggest that strategic controlling of memory engagement is a promising direction for improving both capability and transparency in LMs’ internal thinking behavior.
Limitation.
In this work, we position ourselves as an exploratory study in this area, focusing our evaluation solely on standard SFT. We believe that steered RL could be a more effective approach, as RL may better leverage the capabilities of pretrained LMs compared to SFT in reasoning-intensive settings (Swamy et al., 2025). For scenario like inference-time compute or long CoT, steering with greater utilization on pretrained knowledge may further enhance task performance. Therefore, we also see an opportunity to explore the co-design of algorithms that combine both combine both “internal” steering and “external” CoT generation.
Appendix A Training Details
Hardware and Setup.
All experiments were conducted using a single NVIDIA RTX A6000 GPU with 48 GB of memory. Models with approximately 1 to 2 billion parameters were trained directly on this GPU. For larger models, ranging from 7 to 9 billion parameters, we used DeepSpeed to enable efficient training.
Training Configuration.
All tasks are trained for one epoch. We set the learning rate to 5e-5 and apply a warmup phase of 100 steps. A weight decay of 0.01 is used to regularize the model. We adopt mixed precision training using bfloat16 (bf16) to reduce memory usage and improve efficiency.
The maximum sequence length is 256 tokens. We use a per-device batch size of 4, with gradient checkpointing enabled. To simulate a larger effective batch size and manage memory usage, we use gradient accumulation with 16 steps.
DeepSpeed Optimization.
For large models (7B–9B), we use DeepSpeed with ZeRO Stage 2 optimization. This approach splits the optimizer states and gradients across devices and offloads the optimizer to the CPU. The training gradients are clipped to a maximum norm of 1.0 to stabilize updates. Batch sizes are automatically adjusted by DeepSpeed based on available memory.
Inference Setup.
We use vLLM [Kwon et al., 2023] for all inference runs to ensure efficient memory management and fast decoding. We adopt greedy decoding, and the maximum number of tokens generated per sequence is 256.
Appendix B Fine-Grained Steering during Post-Training
While our intervention and regularization methods provide general approaches to enhance memory vector engagement, finer control mechanisms enable more nuanced steering of parametric knowledge. We present two complementary strategies for achieving fine-grained control: (1) a group-based clustering approach that enables targeted steering of distinct memory regions, and (2) an information surrogate-guided method that selects memory vectors based on their contribution to specific output distributions.
B.1 Group-Based Clustering for Targeted Memory Activation.
B.1.1 Method.
The key coefficients $\mathbf{k}^{(l)}$ can be partitioned into meaningful groups that exhibit distinct activation patterns. Leveraging this structure allows us to apply differentiated steering strategies to various memory regions.
Formally, for each layer $l$, we cluster the key-value pairs $\{(k_{i}^{(l)},v_{i}^{(l)})\}_{i=1}^{d_{m}}$ into $G$ groups $\{\mathcal{G}_{1}^{(l)},\mathcal{G}_{2}^{(l)},\ldots,\mathcal{G}_{G}^{(l)}\}$ based on their functional characteristics. We employ a hierarchical clustering approach with the following steps:
  1. 1.
    Semantic Clustering: First, we cluster memory vectors $v_{i}^{(l)}$ based on their semantic similarity, measured by cosine distance in the embedding space.
  2. 2.
    Activation Pattern Clustering: Second, we sub-cluster based on activation patterns observed during inference on a development set, capturing functional roles within semantic clusters.
For each cluster $\mathcal{G}_{g}^{(l)}$, we define a cluster-specific steering parameter $\beta_{g}$ that modulates the strength of intervention:
$k_{i}^{(l)}\leftarrow k_{i}^{(l)}+\beta_{g}\cdot\Delta k_{i}^{(l)},\text{ for }i\in\mathcal{G}_{g}^{(l)}$ (7)
where $\Delta k_{i}^{(l)}$ is the adjustment magnitude determined by either our intervention or regularization method.
This group-based method offers several advantages. By examining which clusters are active during different tasks, we can selectively strengthen those that contribute most to specific goals, improving task performance. At the same time, we preserve essential clusters that are important for general language understanding, while enhancing those that are underused. Additionally, when aiming to improve factual accuracy, we can prioritize clusters that are associated with factual knowledge.
B.1.2 Results.
Table 6 presents the comprehensive results across all models and benchmarks. Several clear patterns emerge from our experiments:
Memory steering consistently outperforms vanilla SFT. Across all model families and parameter scales, both of our memory clustering approaches demonstrate substantial improvements over standard fine-tuning. The semantic clustering method shows the most dramatic gains, with improvements ranging from 1.2% to 9.7% over vanilla SFT depending on the benchmark and model. These results validate our hypothesis that more balanced memory vector engagement leads to enhanced model capabilities.
Different clustering strategies show task-specific strengths. The semantic clustering approach excels particularly on knowledge-intensive tasks (ARC-c, OBQA) and complex reasoning benchmarks (HellaSwag), achieving improvements of up to +4.7% on ARC-c (Gemma-2-2B) and +3.9% on OBQA (Gemma-2-9B) compared to vanilla SFT. In contrast, activation-based clustering shows more moderate but consistent improvements across a broader range of tasks, suggesting it enables more balanced memory utilization.
Table 6: Performance comparison of Qwen, LLaMA, and Gemma models with different training methods on eight datasets. To highlight improvements, we use blue for significant gains and green for moderate ones. All results are reported as the average scores over three independent runs.
Model Training Accuracy ($\uparrow$) BoolQ PIQA SIQA HellaS. WinoG. ARC-e ARC-c OBQA Small-Scale Models (1-2B parameters) Qwen-2.5-1.5B base model 64.2 78.5 74.3 80.1 76.4 76.9 61.2 75.8      + vanilla SFT 68.5 82.9 79.6 84.8 80.8 81.4 65.8 81.0      + steered SFT w. semantic clustering 70.1 85.2 81.3 94.5 85.3 84.6 69.8 83.2      + steered SFT w. activation clustering 69.4 84.5 80.7 93.2 84.1 83.7 68.5 82.4 LLaMA-3.2-1B base model 65.6 75.3 74.2 78.9 77.8 74.5 60.1 76.3      + vanilla SFT 69.8 79.9 79.5 83.6 82.6 79.8 64.7 81.0      + steered SFT w. semantic clustering 72.6 84.5 81.2 90.3 83.8 85.2 69.4 83.6      + steered SFT w. activation clustering 71.7 83.6 80.4 89.5 83.1 84.1 68.2 82.8 Gemma-2-2B base model 66.5 79.1 73.8 82.7 78.9 77.4 63.8 74.9      + vanilla SFT 70.2 83.4 78.1 87.5 83.3 82.7 68.4 80.1      + steered SFT w. semantic clustering 73.1 86.3 80.8 91.4 87.2 86.5 73.1 85.3      + steered SFT w. activation clustering 72.4 85.7 79.6 90.5 86.3 85.8 72.2 84.1 Large-Scale Models (7-9B parameters) Qwen-2.5-7B base model 68.9 81.2 77.6 87.5 80.3 79.8 65.1 77.6      + vanilla SFT 72.4 84.9 81.5 92.4 84.2 84.2 69.6 82.8      + steered SFT w. semantic clustering 75.2 87.1 82.4 96.2 88.4 87.5 74.8 85.3      + steered SFT w. activation clustering 77.0 86.5 81.9 95.1 87.2 86.3 73.5 84.6 LLaMA-3-8B base model 70.3 85.6 75.7 90.8 81.9 86.2 75.3 80.5      + vanilla SFT 74.6 89.3 79.9 95.5 85.6 90.5 80.4 85.8      + steered SFT w. semantic clustering 78.3 91.0 83.7 96.8 88.3 93.6 82.7 88.4      + steered SFT w. activation clustering 77.2 90.3 82.5 96.1 87.5 92.4 81.8 87.3 Gemma-2-9B base model 71.6 86.3 77.2 90.1 82.5 87.5 77.8 81.7      + vanilla SFT 74.3 90.1 81.7 94.8 86.9 91.7 82.0 86.4      + steered SFT w. semantic clustering 78.5 92.4 84.0 96.9 90.1 94.8 84.5 90.3      + steered SFT w. activation clustering 77.3 91.7 83.2 95.8 89.0 93.6 83.6 89.1
Larger models benefit more significantly from memory steering. While all models show improvements with our methods, the gains are particularly pronounced in the 7-9B parameter models. For instance, semantic clustering improves LLaMA-3-8B by +3.7% on BoolQ and +3.8% on SIQA, compared to more modest gains in the 1B variant. This suggests that larger models contain more untapped parametric knowledge that can be effectively engaged through our steering techniques.
B.2 Information Surrogate-Guided Memory Selection.
Method.
To understand the information introduced by our steering strategy, we extend the key-value formulation from Equation 3 by connecting the FFN output with the final logits:
$\text{FFN}(h)\cdot W_{\text{decode}}=\sum_{i=1}^{d_{m}}k_{i}\cdot(v_{i}\cdot W_{\text{decode}})=\sum_{i=1}^{d_{m}}k_{i}\cdot\phi_{i}$ (8)
where $k_{i}=\sigma(hW_{\text{up}})_{i}$ represents the key coefficient and $\phi_{i}=v_{i}\cdot W_{\text{decode}}$ defines the logit distribution associated with the $i$-th value vector, which we call the information surrogate.
The information surrogate $\phi_{i}$ provides a direct view of how each memory vector influences the final token distribution. This insight enables us to develop a more targeted steering approach:
  1. 1.
    Characterization of Memory Vectors: We analyze the entropy and concentration properties of each $\phi_{i}$ to identify memory vectors that contribute to specific types of generation (e.g., factual statements, reasoning steps, or creative content).
  2. 2.
    Surrogate-Guided Steering: We define a surrogate score function $S(\phi_{i})$ that measures the relevance of each memory vector to our target objective:
    $S(\phi_{i})=\lambda_{1}H(\phi_{i})+\lambda_{2}D_{\text{KL}}(\phi_{i}||\phi_{\text{target}})$ (9)
    where $H(\phi_{i})$ is the entropy of the surrogate distribution, $D_{\text{KL}}$ is the KL divergence from a target distribution $\phi_{\text{target}}$, and $\lambda_{1},\lambda_{2}$ are weighting hyperparameters.
  3. 3.
    Selective Amplification: We modulate key coefficients based on their surrogate scores:
    $k_{i}^{(l)}\leftarrow k_{i}^{(l)}\cdot(1+\gamma\cdot S(\phi_{i}))$ (10)
    where $\gamma$ controls the strength of the surrogate-guided steering.
Theoretical Analysis.
The surrogate-guided selection mechanism introduces several theoretical advantages over conventional fine-tuning approaches. First, we analyze the relationship between surrogate entropy and information capacity. For a memory vector $v_{i}$ with corresponding surrogate $\phi_{i}$, the entropy $H(\phi_{i})$ quantifies the diversity of tokens that can be influenced by this vector. Higher entropy surrogates represent memory vectors that encode distributional knowledge, while low-entropy surrogates often correspond to specialized knowledge concentrated on specific vocabulary subsets. We can formalize this by defining the specificity of a memory vector as $\text{Spec}(v_{i})=1-\frac{H(\phi_{i})}{\log(|V|)}$, where $|V|$ is the vocabulary size. A memory vector with high specificity (low surrogate entropy) exhibits a peaked distribution over the vocabulary, suggesting it encodes precise, specialized information. Conversely, low specificity (high surrogate entropy) indicates a memory vector that contributes more generally across various contexts. This formulation provides a principled approach for analyzing memory vector functionality. Specifically, the information processing capacity of the FFN can be decomposed as:
$\mathcal{I}({\text{FFN}})=\sum{i=1}^{d_{m}}\mathbb{E}_{x\sim\mathcal{D}}[k_{i}(x)]\cdot\text{MI}(\phi_{i};\mathcal{Y})$ (11)
where $\text{MI}(\phi_{i};\mathcal{Y})$ represents the mutual information between the surrogate distribution $\phi_{i}$ and the target next-token distribution $\mathcal{Y}$, and $\mathbb{E}_{x\sim\mathcal{D}}[k_{i}(x)]$ is the expected activation of key $k_{i}$ across the data distribution $\mathcal{D}$. Our surrogate-guided approach can be interpreted as optimizing this information processing capacity by modulating the key coefficients $k_{i}$ according to the information-theoretic properties of their corresponding surrogates $\phi_{i}$. By encouraging the activation of memory vectors with high mutual information with the target distribution, we effectively allocate the model’s capacity toward task-relevant information.
B.3 Results and Analysis
Table 7 presents the performance comparison of our surrogate-guided methods against vanilla SFT across three model families and two size scales. The surrogate-guided selection method consistently outperforms both vanilla SFT and surrogate entropy maximization approaches, with particularly notable gains on reasoning-heavy tasks (BoolQ, SIQA) and knowledge-intensive benchmarks.
Table 7: Performance comparison of information surrogate-based steering. To highlight improvements, we use blue for significant gains and green for moderate ones. All results are reported as the average scores over three independent runs.
Model Training Accuracy ($\uparrow$) BoolQ PIQA SIQA HellaS. WinoG. ARC-e ARC-c OBQA Small-Scale Models (1-2B parameters) Qwen-2.5-1.5B base model 64.2 78.5 74.3 80.1 76.4 76.9 61.2 75.8      + vanilla SFT 68.5 82.9 79.6 84.8 80.8 81.4 65.8 81.0      + steered SFT w. surrogate-guided selection 70.1 84.8 81.7 93.5 85.1 83.6 69.3 79.7      + steered SFT w. surrogate entropy maximization 69.4 84.2 80.2 92.9 84.0 82.9 68.2 78.9 LLaMA-3.2-1B base model 65.6 75.3 74.2 78.9 77.8 74.5 60.1 76.3      + vanilla SFT 69.8 79.9 79.5 83.6 82.6 79.8 64.7 81.0      + steered SFT w. surrogate-guided selection 72.5 84.3 81.6 90.2 83.1 84.5 69.2 82.8      + steered SFT w. surrogate entropy maximization 71.7 83.4 80.7 89.1 82.5 83.1 67.9 82.0 Gemma-2-2B base model 66.5 79.1 73.8 82.7 78.9 77.4 63.8 74.9      + vanilla SFT 70.2 83.4 78.1 87.5 83.3 82.7 68.4 80.1      + steered SFT w. surrogate-guided selection 73.4 86.1 80.6 91.9 86.7 86.1 72.8 84.6      + steered SFT w. surrogate entropy maximization 72.3 85.3 79.2 90.4 85.5 85.3 71.4 83.5 Large-Scale Models (7-9B parameters) Qwen-2.5-7B base model 68.9 81.2 77.6 87.5 80.3 79.8 65.1 77.6      + vanilla SFT 72.4 84.9 81.5 92.4 84.2 84.2 69.6 82.8      + steered SFT w. surrogate-guided selection 75.2 87.3 82.6 95.8 88.5 87.1 74.9 85.3      + steered SFT w. surrogate entropy maximization 76.9 86.2 81.9 94.8 87.0 85.9 73.3 84.0 LLaMA-3-8B base model 70.3 85.6 75.7 90.8 81.9 86.2 75.3 80.5      + vanilla SFT 74.6 89.3 79.9 95.5 85.6 90.5 80.4 85.8      + steered SFT w. surrogate-guided selection 78.3 91.2 83.7 96.8 88.2 93.6 82.4 88.1      + steered SFT w. surrogate entropy maximization 77.1 90.2 81.8 96.0 87.3 92.2 81.5 87.2 Gemma-2-9B base model 71.6 86.3 77.2 90.1 82.5 87.5 77.8 81.7      + vanilla SFT 74.3 90.1 81.7 94.8 86.9 91.7 82.0 86.4      + steered SFT w. surrogate-guided selection 78.6 92.5 84.2 96.9 89.3 94.7 84.8 89.7      + steered SFT w. surrogate entropy maximization 76.9 91.3 83.1 95.5 88.2 93.3 83.5 88.3
Beyond the aggregate statistics, we observe intriguing qualitative differences in how surrogate-guided selection influences model behavior. By examining the top activated memory vectors and their corresponding information surrogates ($\phi_{i}$), we find that our method preferentially engages memory vectors that encode precise factual associations rather than general linguistic patterns. For example, in the ARC-c task, the top-5 memory vectors with the largest positive $\Delta k_{i}$ values predominantly contribute to science concept definitions and physical property relationships. This suggests that surrogate-guided selection effectively identifies and amplifies task-relevant knowledge encoded in specific memory vectors, rather than uniformly increasing memory engagement. Moreover, we find that models trained with surrogate-guided selection exhibit reduced variance in their responses to knowledge-intensive questions, indicating more consistent access to stored parametric knowledge during inference.
Appendix C Understanding Layerwise Contribution
To better understand how different layers contribute to the effectiveness of InfoSteer, we conducted a series of ablation studies. These experiments help us analyze which layers are most sensitive to information steering and which contribute most significantly to overall performance improvements.
C.1 Research Question
We designed our ablation study to investigate the following research questions:
  1. 1.
    Do all layers contribute equally to the information steering effects?
  2. 2.
    Are certain layer groups (early, middle, late) more important for knowledge retrieval?
  3. 3.
    How does the magnitude of steering at different layers affect overall performance?
For these experiments, we used our intervention method with selective application to different layer groups within the model. We also varied the intervention parameters ($p\%$ and $\alpha$) across different layer configurations to understand sensitivity. Table 8 presents the results of our layerwise ablation studies.
Table 8: Layerwise ablation studies. Results show accuracy across different layer configurations with varying intervention parameters.
Model Configuration Layers Avg Acc Base Model 71.4 Vanilla SFT All 72.6      + interv. (early) 1-8 73.8      + interv. (middle) 9-16 75.2      + interv. (late) 17-24 72.9      + interv. (early+middle) 1-16 76.3      + interv. (middle+late) 9-24 74.5      + interv. (all, $\alpha=2$) 1-24 75.5      + interv. (all, $\alpha=5$) 1-24 72.8      + reg. (middle, $\lambda=0.05$) 9-16 74.9      + reg. (all, $\lambda=0.05$) 1-24 74.7      + reg. (all, $\lambda=-0.01$) 1-24 72.3
Layer position matters significantly.
Applying InfoSteer to different layer groups produces notably different results. Early layers (1-8) show moderate improvements, while middle layers (9-16) yield the strongest gains. Late layers (17-24) demonstrate the least improvement, suggesting that information steering is most effective at the intermediate representation level.
Intervention strength sensitivity varies by layer.
Middle layers can tolerate and benefit from stronger interventions ($\alpha=3$), while early layers perform best with moderate intervention ($\alpha=2$), and late layers require gentler steering ($\alpha=1$).
Layer combinations show non-linear effects.
Applying InfoSteer to both early and middle layers yields better-than-additive improvements, suggesting a synergistic effect. However, including late layers tends to diminish these gains, indicating that excessive steering across too many layers may destabilize the model’s representations.
Baseline performance variations.
Our experiments with different regularization strengths ($\lambda$) applied to specific layer groups further confirm that middle layers (9-16) are most receptive to information steering.
These findings demonstrate that information steering should be carefully targeted at specific layers rather than uniformly applied across the entire model. Our optimal configuration focuses on middle and early layers with appropriately calibrated intervention strengths.
Appendix D License for Existing Assets.
Datasets.
The following datasets are used under their respective licenses. For general question answering: BoolQ [Clark et al., 2019] is licensed under CC-BY-SA 3.0, PIQA [Bisk et al., 2019] under the Academic Free License 3.0, SIQA [Sap et al., 2019] and WinoGrande [Sakaguchi et al., 2019] under CC-BY 4.0, HellaSwag [Zellers et al., 2019] under the MIT License, ARC-e and ARC-c [Clark et al., 2018] under CC-BY 4.0, and OBQA [Mihaylov et al., 2018] under the Apache-2.0 License. For arithmetic reasoning: AddSub [Hosseini et al., 2014], MAWPS [Koncel-Kedziorski et al., 2016], MultiArith [Roy and Roth, 2016], and SingleEq [Koncel-Kedziorski et al., 2015] are under CC-BY 4.0, AQuA [Ling et al., 2017] under Apache-2.0, and GSM8K [Cobbe et al., 2021] and SVAMP [Patel et al., 2021] under the MIT License. For instruct-tuning, the Ultrafeedback [Cui et al., 2023] dataset is released under the MIT License.
Model Licenses.
The Qwen-2.5-1.5B and Qwen-2.5-7B models are released under the permissive Apache License 2.0, allowing broad usage including commercial applications [Mesnard et al., 2024]. In contrast, the LLaMA-3.2-1B and LLaMA-3-8B models are distributed under Meta’s custom Llama Community License, which permits research and commercial use but imposes specific restrictions, particularly for organizations with large user bases [Touvron et al., 2023]. The Gemma-2-2B and Gemma-2-9B models are available under Google’s Gemma License, described as commercially friendly; however, access requires users to review and agree to the license terms, typically through platforms like Hugging Face [Mesnard et al., 2024]. Users intending to utilize these models should carefully review the respective licenses to ensure compliance with all terms and conditions.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the paper’s key contributions, which are well-aligned with the theoretical and experimental results presented, accurately reflecting the scope and impact of the work.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss limitations in Section 7 (Conclusion), noting that our approach is exploratory and focused only on standard SFT. We specifically mention that "steered RL could be a more effective approach" and identify opportunities for future work exploring the co-design of algorithms combining "internal" steering with "external" CoT generation.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: Our paper is primarily empirical and doesn’t include formal theoretical results requiring proofs. We build upon the existing key-value memory interpretation of FFN layers, extending it with our steering approach and validating through experiments.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Section 5 provides comprehensive details of our experimental setup, including model architectures (Qwen, LLaMA, Gemma), datasets used for evaluation, and hyperparameters for our steering methods (p= 0.01, $\alpha$ = 1, $\lambda$ = 0.01). Table 2 further details ablation studies with varying hyperparameter settings. We also disclose other details in the appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We release all code for InfoSteer in supplemental material, which will include implementation of both intervention and regularization methods. We use publicly available datasets and models in our experiments.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Section 5.1 details the experimental setup, including base models, datasets, and our default hyperparameters. For all benchmarks, we use the same prompt templates as in prior work. We also report ablation studies in Section 5.3 with different hyperparameter settings.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: As noted in Table 1, all reported results are averaged over three independent runs. Figure 3 clearly displays error bars for our OOD evaluation, and other statistical visualizations (e.g., Figure 4) also include error bars to indicate variance across runs.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We use Nvidia RTX A6000 for training the models.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our research conforms to the NeurIPS Code of Ethics. We use publicly available datasets and models, properly cite all sources, and our method focuses on improving general capabilities rather than applications that could raise ethical concerns.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Our work focuses on improving information utilization in language models, which is fundamental research. While this could potentially lead to more capable AI systems with broader societal implications, our specific technique does not directly introduce new risks beyond those associated with general LM improvements.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our paper introduces a training technique rather than releasing models or datasets that pose risk for misuse. The technique itself is aimed at improving general performance rather than enabling specific capabilities that might require safeguards.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We properly cite all models (Qwen, LLaMA, Gemma) and datasets used in our research. All experiments were conducted in accordance with the terms of use for these assets, using publicly available models and datasets as referenced in Section 5.1.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: Our paper does not introduce new datasets or model weights, but rather a technique for improving existing models.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our research does not involve human subjects or crowdsourcing.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our research does not involve human subjects, so IRB approval was not required.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: While our paper studies LLMs, we did not use LLMs in the development of our methodology or for content generation during the research process.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.